# TrustRate: A Decentralized Platform for Hijack-Resistant Anonymous Reviews

### v1

@@ 2.2.2. Hijack Resistance
Another scenario in this system is that every honest user, on average, participates only in some small number, say, $c$ reviews in time $T$ (which captures common review systems today). Informally, if all polls have an equal number of votes from honest users, then each poll has $\kappa\approx\frac{nc}{M}$ honest reviews This assumption can be easily removed probabilistically. Now, if $\rho=0.5$, an adversary would only need to control $\kappa$ users to hijack all polls. This system is hijack resistant with very poor parameters – namely, $\gamma\approx\frac{\kappa}{n}$ (for $\theta=1$, $\rho=0.5$) More generally, $\gamma\approx\frac{\rho}{1-\rho}\cdot\frac{\kappa}{n}$..
@@ 4.3. Voting on a Poll
To overcome this attack of multiple votes by the same user, we make a few changes to the $\mathsf{URS}$ protocol (from §3.2). In particular, in the original scheme, the tag $\tau$ is generated as a function of the message $M$ that does not suffice to prevent double voting. We change the scheme to have an additional tag $\nu$ that only depends on the poll ID We cannot use only the tag that depends on PID since we need to sign the vote value as well.:
@@ 4.4. Dynamic Random Access Control (DRAC)
where $pk_{u}$ is the public key of user $u$ and $B_{wait}$ is a constant parameterThe parameter $B_{wait}$ determines the number of blocks users need to wait after a poll is committed before being able to calculate the $\mathsf{VRF}$s and determine the eligibility to vote. As soon as $\textsf{seed}_{i+B_{wait}}$ is available, the set of users eligible to vote on the poll, $\textsf{Voters}^{(P)}$, can be computed by anyone. See Lemma G.1 for details on how $B_{wait}$ is calculated to be $38$ blocks. Based on this $\mathsf{VRF}$, we select the n_req highest $\mathsf{VRF}$s to be the reviewers for this poll, i.e., $\textsf{Voters}^{(P)}$.
@@ 4.6. Threat Model
In our threat model, similar to Blockene, we assume that atleast $20\%$ of Politicians and $75\%$ of all Citizens are honest. The remaining $80\%$ of Politicians and $25\%$ of Citizens can be malicious and can collude with each other. One distinction to note is that malicious users in Blockene Citizens and malicious users in the definition of hijacking cost (Defn. 2.1) do not necessarily refer to the same set of users or the same set of behaviours Malicious behaviour in Citizens and Politicians primarily involves attacks on liveness or soundness of the block commit protocol. Malicious behaviour in the TrustRate polling system involves skewing the votes on the polls themselves to paint an agenda or promote a specific cause. . In TrustRate’s permissionless use case, every Citizen is uniquely linked to a $\mathsf{URS}_{m}$ key and both these sets of users are the same. However, in the permissioned setting, the users registering to vote on TrustRate and the Citizens could possibly be differing sets of users.
@@ 5.1.2. Verifying ring signatures
A Citizen queries a safe sample of Politicians to check whether the signatures on all the transactions being considered are valid. If not, the Politicians return a list of invalid transactions that are then explicitly verified by the Citizen. In the common case, when a bulk of the transactions are valid, most of the verification gets off-loaded. If a Politician tries to slow down a Citizen by claiming correct transactions to be incorrect the Politician can be blacklisted.All messages exchanged between nodes are authenticated by the sender. Hence, an incorrect claim by a Politician can easily be used to blacklist them.
Note that Citizens do not even have to obtain the ring to verify most $\mathsf{URS}$ vote transactions – they can directly ask a safe sample of Politicians whether the signatures in a set of transactions are all correct. If any transaction is reported as incorrect, Citizens can ask a Politician for the ring of eligible voters for that poll – that Politician then provides a list of voters, along with proof.Recall that for every new poll, the block committee commits the hash of the ring, $H(R)$, to the global state. Thus, any Citizen can query the global state (with Merkle Tree challenge path) to verify if the ring is correct. (additional details: Appendix F.2).
@@ 5.1.3. Batch verification of signatures
In Blockene, the decision on which transactions to include in the current block is made by a deterministic, random subset of $45$ Politicians every block. Each of these selected Politicians commit to a $\mathsf{tx\_pool}$, a set of transactions UUIDs that they intend to include in the current block. To ensure that multiple Politicians don’t commit to the same set of transactions, every transaction is also uniquely, deterministically mappedIn round $i$, allow only the $j^{th}$ Politician out of these $45$ to include a certain vote $\mathcal{V}$ in it’s $\mathsf{tx\_pool}$, where $j\equiv\mathsf{Hash}(i||\mathcal{V})\bmod 45$ to one of these $45$ politicians, essentially sharding the pending transactions into disjoint sets. In such a system, even if a Politician had multiple votes from the same poll in its pool of waiting transactions, it would not be able to group and commit to all of them, since some votes might be in a different shard.
@@ 6.2. Experimental Setup
We simulate a TrustRate instance with one million users who participate by sending in votes when selected. By default, we use a ring size of $100$, i.e., hundred TrustRate users are selected to vote in every pollPerformance numbers for other ring sizes are in Section 6.3.5. We continuously create and bombard the system with enough new polls (each of which is $\approx 0.35$ KB) to ensure that Politicians’ pending transaction pools are always sufficiently full. Each review consists of a 256-byte field - sufficient for both a numeric rating and a short text response, along with a signature of the reviewer. The signature size depends on the size of the ring being used (e.g., with ring size 100, signature size is $\approx$ 1.95KB and the total size of content + signature is 2.2KB).
@@ 6.3.3. Impact of batching
We perform microbenchmarks on Politicians to demonstrate the impact batch verification of $\mathsf{URS}_{m}$ signatures can haveNote that all other results present E2E throughputs for the most pessimistic case where every vote needs to be individually verified with no batching.. Table 4 presents results on time taken to batch-verify votes in rings ($\mathcal{R}$) of various sizes. From the table, we see that verifying even two signatures at once provides gains of $1.3$x at $\mathcal{R}=50$ to $1.8$x at $\mathcal{R}=1000$. For $\mathcal{R}=100$, each block has ${\approx}4000$ votes in it, and would take ${\approx}66$s to verify serially; batching would reduce this to $31.6$s with batch size 2 or $23.6$s with batch size 8.
@@ 6.3.4. Cost of anonymity
In the previous section, we examined how our modifications to the Blockene protocol (and our design of $\mathsf{URS}_{m}$’s batch-verify) has improved the performance of TrustRate by a significant factor. In this section, we attempt to quantify the additional cost of adding $\mathsf{URS}_{m}$ and blind signatures. Row M1 of Table 5 is a baseline scenarioNote that the 262.75 votes/sec indicated for scenarios M1 and M2 includes usage of the ring offload optimisation we developed in §5.1.1. Without this optimization, using a non-modified Blockene backend would result in a throughput of 23.8 votes/sec instead - more than 10×10$\times$10 × lesser. where there is no anonymity at all - users register to the system using a pseudonymous EdDSA key, and use that key for all interactions with the underlying Blockene protocol, as well as for casting votes on polls they are selected for. This key is linked either to their TEE (in the public use case) or to their employee ID (in the organisational case). M2 is a similar scenario, which can only be used for the permissioned organisational use-case, where blind signatures (§4.1.1) are used to ensure EdDSA keys are not linkable to employee IDs.
@@
We notice from Table 5 that the system performance in (M1, M2) and (M3, M4) are the same – adding membership anonymity requires a change only in the setup phase. Once the blind signature protocol has been executed, the steady state throughput of the system depends on the voting scheme alone. Appendix A.1 contains microbenchmarks that quantify the cost of this one-time setup phase. For some organisational settings, using a TrustRate-like system with only membership anonymity might be a suitable option, since it is 5×5$\times$5 × fasterThe only downside in not having content anonymity in a scenario like this is that adversaries will be able to build voting profiles of users, by compiling a list of polls each key voted on, and the corresponding value of that vote. Using $\mathsf{URS}_{m}$ would prevent attacks such as these..
@@ 6.3.6. Participation costs
We profiled and collected data on resource usage by participants of TrustRate in the fully honest case of §6.3.1. In this scenario, around $45,000$ new polls are created per day, and every user on average takes part in $\approx 5$ pollsSee Appendix F.4.1 for details.. Similar to Blockene, the data usage in TrustRate was observed to be 20MB/block committee and the cost of periodic GetLedgers (§3.1) was $21$MB/day. With an observed block latency of $\approx 72$ seconds, corresponding to 1200 blocks/day, each user on average participates in $\approx\frac{2000\times 1200}{10^{6}}=2.4$ block committees per day (i.e., $40-60$MB/day total). Thus, the total network usage for a Citizen in our system is $70$MB/day on average – only a slight increase over Blockene’s $60$MB/day usage.
@@ Appendix D Handling User Apathy and Unavailability
In the rest of this paper we stick to the simpler definitions of threshold updates (in Equations 8, 9) in lieu of these normalized forms for simplicity.One problem with this dynamic system is that of permanently inactive subscribers. If someone subscribes to a topic, and then just uninstalls the app or never participates in the system again, the thresholding system will still ensure that enough actual votes are recorded on each poll. However, this comes at an additional cost of a bigger ring size which might translate to longer signatures and higher sign/verify times (Appendix B.1). This, however, can be easily solved with a resubscription mandate: every user must resubscribe to topics after a certain period of time (e.g., a month).
@@ Appendix E Building TrustRate  over Blockene - Details
Attribute
Notation
Size (B)
Explanation
@@
SHA256
$H(...)$
32
SHA256 hash of content in brackets
Block num.
$B_{i}$
8
Block $i$.
@@
Poll ID
PID
8
Unique identifier for each poll.
Topic ID
$\textsf{Topic}_{j}$
4
Refers to the $j^{th}$ topic
Poll Data
PollContent
256
The question being asked in the poll.
Req. votes
n_req
4
Number of votes required on a poll
Seen votes
n_seen
4
Number of votes received on a poll so far
Vote Window
$B_{vw}$
4
Number of blocks for which poll is open for voting.
@@
Vote value
$\mathcal{V}$
257
1B rating field (eg. 5-star scale) + 256B free-form text component (eg. review) of size 256-bytes
$\mathsf{URS}_{m}$ tag
$\nu$
32
$\mathsf{URS}_{m}$ tags ensure uniqueness of votes
@@
Votes seen
$v_{seen}$
4
Total votes received across all polls in a certain topic, in current epoch
@@ 6 ScalingFactor ”U”, $\textsf{Topic}_{j}$ $v_{exp},v_{seen}$ 5+8=10
    • Poll contains essential information or meta data about the poll (i.e. information that is needed to verify a new vote). This includes the block number in which the poll was committed ($B_{n}$), number of votes seen so far (n_seen), the $B_{vw}$ for the poll and a succinct representation for eligible voters $H(R)$ where $R$ is the ring to be used in ring signatures during votingSee Section 6.3.6 for a description on how users identify if they are part of the poll. When a poll is committed, $R$ is unknown and only gets updated after $B_{wait}$ number of blocks as discussed in Section E.4.
@@ F.1. Ring Calculation Offload
[…]
The possible malicious behaviours a Politician could use to attack this protocol are: (1) provide Citizens with the incorrect ring-hash, or (2) refuse to respond to requests for the ring. The first form of attack is an attack on the correctness of our protocol while the latter is an attack on the liveness. Lemma F.1 discusses the soundness of Algorithm 1. To prevent a single malicious node from causing an indefinite stall of the block committee in this step, we set a timeout ($\mathcal{T}_{h}$) on the request based on a calculation of the maximum conceivable time it could take for a honest Politician to respond. This calculation of $\mathcal{T}_{h}$ has two parameters - $t_{\mathsf{hash}}$ and $n_{\mathsf{thread}}$. The first parameter is the average time taken by a Politician to perform one $\mathsf{Hash}$ computationFor SHA256 hashes on our Politician nodes, we observed that $10^{6}$ hashes take $\approx 1$ second on a single core. and the second is the minimum number of parallel threads a Politician is expected to have for performing hash computations. Both of these are parameters that can be determined before the system starts via microbenchmarks and analysing the system configuration. In our experiments (§6), we noticed that honest Politicians are able to complete all required hash computations for rings well in advance of when they actually become required (after GSRead).
Lemma F.0 ().
@@ G.2. Proofs for fairness and anonymity
This follows from the fact that the opinion committee is selected at random from the set of all eligible users (guaranteed by $\mathsf{VRF}$) and from Lemma G.2 (which ensures that every eligible user can only cast one vote per poll). For some small probability $\epsilon$, the guarantee holds with $\delta_{f}$ given by the Chernoff bound (Lemma G.5) due to the random sampling.
Lemma G.0 ().

### v3

@@ 2.2.2. Hijack Resistance
Another scenario in this system is that every honest user, on average, participates only in some small number, say, $c$ reviews in time $T$ (which captures common review systems today). Informally, if all polls have an equal number of votes from honest users, then each poll has $\kappa\approx\frac{nc}{M}$ honest reviews  [^1] This assumption can be easily removed probabilistically. Now, if $\rho=0.5$, an adversary would only need to control $\kappa$ users to hijack all polls. This system is hijack resistant with very poor parameters – namely, $\gamma\approx\frac{\kappa}{n}$ (for $\theta=1$, $\rho=0.5$)  [^2] More generally, $\gamma\approx\frac{\rho}{1-\rho}\cdot\frac{\kappa}{n}$..
@@ 4.3. Voting on a Poll
To overcome this attack of multiple votes by the same user, we make a few changes to the $\mathsf{URS}$ protocol (from §3.2). In particular, in the original scheme, the tag $\tau$ is generated as a function of the message $M$ that does not suffice to prevent double voting. We change the scheme to have an additional tag $\nu$ that only depends on the poll ID  [^3] We cannot use only the tag that depends on PID since we need to sign the vote value as well.:
@@ 4.4. Dynamic Random Access Control (DRAC)
where $pk_{u}$ is the public key of user $u$ and $B_{wait}$ is a constant parameter [^4] The parameter $B_{wait}$ determines the number of blocks users need to wait after a poll is committed before being able to calculate the $\mathsf{VRF}$s and determine the eligibility to vote. As soon as $\textsf{seed}_{i+B_{wait}}$ is available, the set of users eligible to vote on the poll, $\textsf{Voters}^{(P)}$, can be computed by anyone. See Lemma G.1 for details on how $B_{wait}$ is calculated to be $38$ blocks. Based on this $\mathsf{VRF}$, we select the n_req highest $\mathsf{VRF}$s to be the reviewers for this poll, i.e., $\textsf{Voters}^{(P)}$.
@@ 4.6. Threat Model
In our threat model, similar to Blockene, we assume that atleast $20\%$ of Politicians and $75\%$ of all Citizens are honest. The remaining $80\%$ of Politicians and $25\%$ of Citizens can be malicious and can collude with each other. One distinction to note is that malicious users in Blockene Citizens and malicious users in the definition of hijacking cost (Defn. 2.1) do not necessarily refer to the same set of users or the same set of behaviours 555 Malicious behaviour in Citizens and Politicians primarily involves attacks on liveness or soundness of the block commit protocol. Malicious behaviour in the TrustRate polling system involves skewing the votes on the polls themselves to paint an agenda or promote a specific cause. . In TrustRate’s permissionless use case, every Citizen is uniquely linked to a $\mathsf{URS}_{m}$ key and both these sets of users are the same. However, in the permissioned setting, the users registering to vote on TrustRate and the Citizens could possibly be differing sets of users.
@@ 5.1.2. Verifying ring signatures
A Citizen queries a safe sample of Politicians to check whether the signatures on all the transactions being considered are valid. If not, the Politicians return a list of invalid transactions that are then explicitly verified by the Citizen. In the common case, when a bulk of the transactions are valid, most of the verification gets off-loaded. If a Politician tries to slow down a Citizen by claiming correct transactions to be incorrect the Politician can be blacklisted. [^7] All messages exchanged between nodes are authenticated by the sender. Hence, an incorrect claim by a Politician can easily be used to blacklist them.
Note that Citizens do not even have to obtain the ring to verify most $\mathsf{URS}$ vote transactions – they can directly ask a safe sample of Politicians whether the signatures in a set of transactions are all correct. If any transaction is reported as incorrect, Citizens can ask a Politician for the ring of eligible voters for that poll – that Politician then provides a list of voters, along with proof. [^8] Recall that for every new poll, the block committee commits the hash of the ring, $H(R)$, to the global state. Thus, any Citizen can query the global state (with Merkle Tree challenge path) to verify if the ring is correct. (additional details: Appendix F.2).
@@ 5.1.3. Batch verification of signatures
In Blockene, the decision on which transactions to include in the current block is made by a deterministic, random subset of $45$ Politicians every block. Each of these selected Politicians commit to a $\mathsf{tx\_pool}$, a set of transactions UUIDs that they intend to include in the current block. To ensure that multiple Politicians don’t commit to the same set of transactions, every transaction is also uniquely, deterministically mapped [^9] In round $i$, allow only the $j^{th}$ Politician out of these $45$ to include a certain vote $\mathcal{V}$ in it’s $\mathsf{tx\_pool}$, where $j\equiv\mathsf{Hash}(i||\mathcal{V})\bmod 45$ to one of these $45$ politicians, essentially sharding the pending transactions into disjoint sets. In such a system, even if a Politician had multiple votes from the same poll in its pool of waiting transactions, it would not be able to group and commit to all of them, since some votes might be in a different shard.
@@ 6.2. Experimental Setup
We simulate a TrustRate instance with one million users who participate by sending in votes when selected. By default, we use a ring size of $100$, i.e., hundred TrustRate users are selected to vote in every poll [^10] Performance numbers for other ring sizes are in Section 6.3.5. We continuously create and bombard the system with enough new polls (each of which is $\approx 0.35$ KB) to ensure that Politicians’ pending transaction pools are always sufficiently full. Each review consists of a 256-byte field - sufficient for both a numeric rating and a short text response, along with a signature of the reviewer. The signature size depends on the size of the ring being used (e.g., with ring size 100, signature size is $\approx$ 1.95KB and the total size of content + signature is 2.2KB).
@@ 6.3.3. Impact of batching
We perform microbenchmarks on Politicians to demonstrate the impact batch verification of $\mathsf{URS}_{m}$ signatures can have [^11] Note that all other results present E2E throughputs for the most pessimistic case where every vote needs to be individually verified with no batching.. Table 4 presents results on time taken to batch-verify votes in rings ($\mathcal{R}$) of various sizes. From the table, we see that verifying even two signatures at once provides gains of $1.3$x at $\mathcal{R}=50$ to $1.8$x at $\mathcal{R}=1000$. For $\mathcal{R}=100$, each block has ${\approx}4000$ votes in it, and would take ${\approx}66$s to verify serially; batching would reduce this to $31.6$s with batch size 2 or $23.6$s with batch size 8.
@@ 6.3.4. Cost of anonymity
In the previous section, we examined how our modifications to the Blockene protocol (and our design of $\mathsf{URS}_{m}$’s batch-verify) has improved the performance of TrustRate by a significant factor. In this section, we attempt to quantify the additional cost of adding $\mathsf{URS}_{m}$ and blind signatures. Row M1 of Table 5 is a baseline scenario [^12] Note that the 262.75 votes/sec indicated for scenarios M1 and M2 includes usage of the ring offload optimisation we developed in §5.1.1. Without this optimization, using a non-modified Blockene backend would result in a throughput of 23.8 votes/sec instead - more than 10×10$\times$10 × lesser. where there is no anonymity at all - users register to the system using a pseudonymous EdDSA key, and use that key for all interactions with the underlying Blockene protocol, as well as for casting votes on polls they are selected for. This key is linked either to their TEE (in the public use case) or to their employee ID (in the organisational case). M2 is a similar scenario, which can only be used for the permissioned organisational use-case, where blind signatures (§4.1.1) are used to ensure EdDSA keys are not linkable to employee IDs.
@@
We notice from Table 5 that the system performance in (M1, M2) and (M3, M4) are the same – adding membership anonymity requires a change only in the setup phase. Once the blind signature protocol has been executed, the steady state throughput of the system depends on the voting scheme alone. Appendix A.1 contains microbenchmarks that quantify the cost of this one-time setup phase. For some organisational settings, using a TrustRate-like system with only membership anonymity might be a suitable option, since it is 5×5$\times$5 × faster [^13] The only downside in not having content anonymity in a scenario like this is that adversaries will be able to build voting profiles of users, by compiling a list of polls each key voted on, and the corresponding value of that vote. Using $\mathsf{URS}_{m}$ would prevent attacks such as these..
@@ 6.3.6. Participation costs
We profiled and collected data on resource usage by participants of TrustRate in the fully honest case of §6.3.1. In this scenario, around $45,000$ new polls are created per day, and every user on average takes part in $\approx 5$ polls [^14] See Appendix F.4.1 for details.. Similar to Blockene, the data usage in TrustRate was observed to be 20MB/block committee and the cost of periodic GetLedgers (§3.1) was $21$MB/day. With an observed block latency of $\approx 72$ seconds, corresponding to 1200 blocks/day, each user on average participates in $\approx\frac{2000\times 1200}{10^{6}}=2.4$ block committees per day (i.e., $40-60$MB/day total). Thus, the total network usage for a Citizen in our system is $70$MB/day on average – only a slight increase over Blockene’s $60$MB/day usage.
@@ Appendix D Handling User Apathy and Unavailability
In the rest of this paper we stick to the simpler definitions of threshold updates (in Equations 8, 9) in lieu of these normalized forms for simplicity. [^15] One problem with this dynamic system is that of permanently inactive subscribers. If someone subscribes to a topic, and then just uninstalls the app or never participates in the system again, the thresholding system will still ensure that enough actual votes are recorded on each poll. However, this comes at an additional cost of a bigger ring size which might translate to longer signatures and higher sign/verify times (Appendix B.1). This, however, can be easily solved with a resubscription mandate: every user must resubscribe to topics after a certain period of time (e.g., a month).
@@ Appendix E Building TrustRate  over Blockene - Details
Attribute Notation Size (B) Explanation
@@
SHA256 $H(...)$ 32 SHA256 hash of content in brackets
Block num. $B_{i}$ 8 Block $i$.
@@
Poll ID PID 8 Unique identifier for each poll.
Topic ID $\textsf{Topic}_{j}$ 4 Refers to the $j^{th}$ topic
Poll Data PollContent 256 The question being asked in the poll.
Req. votes n_req 4 Number of votes required on a poll
Seen votes n_seen 4 Number of votes received on a poll so far
Vote Window $B_{vw}$ 4 Number of blocks for which poll is open for voting.
@@
Vote value $\mathcal{V}$ 257 1B rating field (eg. 5-star scale) + 256B free-form text component (eg. review) of size 256-bytes
$\mathsf{URS}_{m}$ tag $\nu$ 32 $\mathsf{URS}_{m}$ tags ensure uniqueness of votes
@@
Votes seen $v_{seen}$ 4 Total votes received across all polls in a certain topic, in current epoch
@@ 6 ScalingFactor ”U”, $\textsf{Topic}_{j}$ $v_{exp},v_{seen}$ 5+8=10
    • Poll contains essential information or meta data about the poll (i.e. information that is needed to verify a new vote). This includes the block number in which the poll was committed ($B_{n}$), number of votes seen so far (n_seen), the $B_{vw}$ for the poll and a succinct representation for eligible voters $H(R)$ where $R$ is the ring to be used in ring signatures during voting [^16] See Section 6.3.6 for a description on how users identify if they are part of the poll. When a poll is committed, $R$ is unknown and only gets updated after $B_{wait}$ number of blocks as discussed in Section E.4.
@@ F.1. Ring Calculation Offload
Algorithm 1 describes the ring offload protocol. In this protocol, Citizens are attempting to verifiably learn the ring-hash, a hash of list of eligible $\textsf{Voters}^{(P)}_{j}$ for a certain new poll Pj. In the best case, when all Politicians respond with the same value, the total network cost of this protocol is $(32\cdot m\textrm{ bytes})=0.8$KB per poll, where $m=25$ is the safe sample size, and we assume the output of the Hash function is 256 bits (or 32-bytes).
The possible malicious behaviours a Politician could use to attack this protocol are: (1) provide Citizens with the incorrect ring-hash, or (2) refuse to respond to requests for the ring. The first form of attack is an attack on the correctness of our protocol while the latter is an attack on the liveness. Lemma F.1 discusses the soundness of Algorithm 1. To prevent a single malicious node from causing an indefinite stall of the block committee in this step, we set a timeout ($\mathcal{T}_{h}$) on the request based on a calculation of the maximum conceivable time it could take for a honest Politician to respond. This calculation of $\mathcal{T}_{h}$ has two parameters - $t_{\mathsf{hash}}$ and $n_{\mathsf{thread}}$. The first parameter is the average time taken by a Politician to perform one $\mathsf{Hash}$ computation [^17] For SHA256 hashes on our Politician nodes, we observed that $10^{6}$ hashes take $\approx 1$ second on a single core. and the second is the minimum number of parallel threads a Politician is expected to have for performing hash computations. Both of these are parameters that can be determined before the system starts via microbenchmarks and analysing the system configuration. In our experiments (§6), we noticed that honest Politicians are able to complete all required hash computations for rings well in advance of when they actually become required (after GSRead).
Lemma F.0.
@@ G.2. Proofs for fairness and anonymity
This follows from the fact that the opinion committee is selected at random from the set of all eligible users (guaranteed by $\mathsf{VRF}$) and from Lemma G.2 (which ensures that every eligible user can only cast one vote per poll). For some small probability $\epsilon$, the guarantee holds with $\delta_{f}$ given by the Chernoff bound (Lemma G.5) due to the random sampling. ∎
Lemma G.0.